### v1

@@
Abstract
Abstract: Vineyards are a common way to study persistence diagrams of a data set which is changing, as strong stability means that it is possible to pair points in “nearby” persistence diagrams, yielding a family of point sets which connect into curves when stacked. Recent work has also studied monodromy in the persistent homology transform, demonstrating some interesting connections between an input shape and monodromy in the persistent homology transform for 0-dimensional homology embedded in $\mathbb{R}^{2}$. In this work, we re-characterize monodromy in terms of periodicity of the associated vineyard of persistence diagrams.
We construct a family of objects in any dimension which have non-trivial monodromy for $l$-persistence of any periodicity and for any $l$. More generally we prove that any knot or link can appear as a vineyard for a shape in $\mathbb{R}^{d}$, with $d\geq 3$. This shows an intriguing and, to the best of our knowledge, previously unknown connection between knots and persistence vineyards. In particular this shows that vineyards are topologically as rich as one could possibly hope.
keywords:
Monodromy, Knots and links, Persistent homology, Vineyards
@@ 1 Introduction
Computational topology has a number of different branches which, despite some overlap, have remained fairly distinct in nature. On the one hand there is persistent homology [18, 22, 16], a more recent and active area which has its roots in algebraic topology; see [21, 14] for recent survey books on this active topic. On the other hand there are the computational aspects of knot theory and the study and characterization of low dimensional manifolds’ topology, which has a long history of algorithmic development, perhaps dating back to Dehn’s algorithm [12, 13] and the many following algorithmic results in more recent decades on shape and knot recognition in low dimensions [23, 27, 19, 5, 6].
@@
The first goal of this paper is studying an interesting new link between the these two important branches of computational topology. Thanks to the work of Alexander [2] we know that every knot or link can be represented as a braid, that is, for every link there is a braid such that if we glue the ends of the braid, we recover the link. Recall also that for a continuous one parameter family of filtrations, we can “stack” the persistence diagrams of these filtrations; we call the resulting object a vineyard [11]. Thanks to the stability of persistence diagrams, the points in the persistence diagram move continuously (even Lipschitz continuously) with the parameter. This means that we can follow a point in (the stack of) the persistence diagrams; the resulting curve is called a vine. We will prove that for every link there exists an embedded manifold and a family of functions on $\mathcal{M}$ (where each function is induced by the distance to a point in the ambient space, and where in turn each point comes from a curve $\gamma$) such that the vineyard of the family of function yields the braid representing the knot or link in the sense of Alexander.
@@
The second goal of this paper is to show that any type of monodromy can occur in vineyards. This is part of a new research direction in computational topology. In [3], the occurrence of monodromy in the context of the directional persistence transform is studied, more precisely for 0-dimensional persistence modules of objects embedded in $\mathbb{\mathbb{R}}^{2}$. They conclude with an open question about demonstrating monodromy in higher dimensions, as well as several interesting and more open ended questions related to better understanding what monodromy is capturing about the input shape.
1.1 Our contributions
In this paper we exhibit that any type of monodromy and the braid associated to any knot or link can occur in a vineyard. To make our statement more precise, however, we need to introduce some nomenclature, although full definitions will be deferred until Section 2.
@@ 1.1 Our contributions
Monodromy is the effect where if one makes a loop in a base space of a covering or fibre bundle, the lifted curve may not end up in the same point as you started out with. We say the monodromy is of period $2\pi k$ (with $k>0$) if the lifted curve returns to starting point after $k$ revolutions in the base space. In our context the base space is a closed curve or loop $\gamma:[0,2\pi]\to\mathbb{R}^{d}$ into which we have embedded a manifold $\mathcal{M}$ (mostly some knot, link or some offset of a knot or link). The fibres are the persistence diagrams of the distance function restricted to the manifold $\mathcal{M}$, that is $d(x,\gamma(t))_{\mathcal{M}}$. The bundle therefore is the vineyard. The lifted curve is a vine $\tilde{\gamma}_{(b_{0},d_{0})}(t)$ in the vineyard starting at $(b_{0},d_{0})$ in the persistence diagram of $d(x,\gamma(0))_{\mathcal{M}}$ and the periodicity is the smallest $k>0$, such that for all $i$, $\tilde{\gamma}_{(b_{0},d_{0})}(0)=\tilde{\gamma}_{(b_{i},d_{i})}(2\pi k)$, where we assume that the vine is non-degenerate in the sense that it stays away from the diagonal.
@@
Theorem 1.1.
The persistence distance transform in $\mathbb{R}^{d}$ can exhibit monodromy for persistence up to the ($d-2$)th homology and for extended persistence up to the ($d-1$)th homology. Moreover the periodicity of the monodromy can be $2k\pi$ for any $k\in\mathbb{Z}_{\geq 2}$.
Figure 1: For a given a link (top), consisting of a trefoil and a circle, we construct an angular braid embedded near an annulus (bottom left), which we then modify (bottom right) such that vineyard consisting of the persistence diagrams of the distance function to a point following the twisted annulus on the ‘outside’ contains the input braid with some surgery. See also Figure 2 and the corresponding Example in Appendix A.
Figure 2: Top: side view of the embedding of the link depicted in Figure 1 following a twisted annulus, see also the proof of Theorem 1.1 and the corresponding Example in Appendix A. The observation loop is shown in light blue, with the front part darkened to correspond with the period shown in the bottom figure. Middle: Front-angled view of the embedded link. Bottom: the fraction of the full period (from 00 to $2\pi$) depicted in dark blue above of the braided vineyard of this link, which captures all crossing in the vineyard and exhibits monodromy of period $2\pi\cdot 3$. To transform this braided vineyard into a closed braid, identify the “sides”, or slices, at 00 and $2\pi$.
Theorem 1.2.
Given a knot or link, $d,l\in\mathbb{Z}_{>0}$, with $d\geq 3$ and $l<d-2$, then there exists an $\mathcal{M}\subset\mathbb{R}^{d}$ and a closed curve $\gamma\subset\mathbb{R}^{d}$ such that by identifying the ends of the $l$-vineyard of $d(x,\gamma(t))_{\mathcal{M}}$ will yield a knot or link, which contains the given knot or link as a subset, that is, it is topologically the knot or link we started our construction with after removing some spurious connected components.
It is worth emphasizing that in our constructive proof, the choice of not only $\mathcal{M}$ but also $\gamma$ is critical, as there are choices of $\gamma$ which yield no links in the vineyard.
@@ 2.1 Monodromy
Monodromy is an important concept in mathematics that appears in various guises. We refer to the review [15] (the first part of which is almost a review of reviews) and the other reviews mentioned in that paper for an overview of the various aspects of the theory. In this paper we will only consider the simplest incarnation, and only in the setting of topology data analysis.
Let $\tilde{X}$ be a covering space of $X$ with covering map $C:\tilde{X}\to X$, that is for every $x\in X$ there exists an open neighbourhood $x\in U$ and a discrete set $J$, such that $C^{-1}(U)=\sqcup_{i\in J}V_{i}$ and $C|_{V_{j}}:V_{j}\to U$ is a homeomorphism for all $j\in J$. We call the inverse images of points $x\in X$ of the map $C$ the fibres. For a curve $\gamma:[0,2\pi]\to X$ we write $\tilde{\gamma}$ for (one of) its lift(s), that is a continuous map $\tilde{\gamma}:[0,2\pi]\to\tilde{X}$ such that $C\cdot\tilde{\gamma}=\gamma$. If $\gamma$ is a loop, that is $\gamma(0)=\gamma(2\pi)$, then we say that $\gamma$ exhibits monodromy (at the starting point $\tilde{\gamma}(0)\in C^{-1}(\gamma(0))$) if we have that the start and end points of its lift $\tilde{\gamma}$ are different, i.e. $\tilde{\gamma}(0)\neq\tilde{\gamma}(2\pi)$. The difference between $\tilde{\gamma}(0)$ and $\tilde{\gamma}(2\pi)$ is also referred to as monodromy (this difference can in certain cases be best represented by a group, see [15], although we will not need this in our discussion).
Figure 3: Here we see a cover $\tilde{X}$ (in this case a double cover) of the base space $X$, in this case a circle as well as the curve $\gamma$ and its lift $\tilde{\gamma}$.
Because $\gamma$ is a loop we can extend it, formally speaking by concatenating with itself. Here we adopt the convention that the if $\gamma$ and $\tilde{\gamma}$ are two curves parametrized by $[0,2\pi]$, then the concatenation $\gamma\circ\tilde{\gamma}$ is parametrized by $[0,4\pi]$, that is we don’t rescale the parametrization interval. We write
@@
where $\tilde{\gamma}$ is the lifting of $\gamma$. In fact the right hand side does not even have to be a continuous curve. We say that a loop $\gamma$ (parametrized by $[0,2\pi]$) in the base space $X$ exhibits monodromy of order $k$ if $k$ is the smallest positive integer such that the lifted curve $\widetilde{\gamma^{k}}$ satisfies
@@ 2.2 Knots, links, and braids
In this subsection we briefly recall the formal definitions of knots, links and braids. An (oriented) knot is the the equivalence class of oriented closed curves embedded in $3$-dimensional Euclidean space, $\gamma\colon\mathbb{S}^{1}\to\mathbb{R}^{3}$, under ambient isotopy. To simplify the discussion, we will abuse notation and write $\gamma$ for the map as well as its image in $\mathbb{R}^{3}$. A link with $n$ components is a disjoint union of $n$ knots, $L=\gamma_{1},\cup\ldots\cup\gamma_{n}\subset\mathbb{R}^{3}$. We say two knots (or links) are equivalent if there is an orientation-preserving homeomorphism from $\mathbb{R}^{3}$ to $\mathbb{R}^{3}$ such that one knot (or link) is the image of the other.
For each $u\in\mathbb{S}^{2}$, the projection of a knot (or link), $\gamma$, in the direction $u$ provides a knot diagram (or link diagram) of $\gamma$ with crossings. In this paper we will assume diagrams with generic crossings.
A braid on $m$ strands is the disjoint union of $m$ intervals embedded in a solid cylinder, $B=B_{i}\colon I\to D^{2}\times I$, monotonically increasing with respect to $I$, such that $B_{i}(0)=(d_{i},0)$ and $B_{i}(1)=(d_{j},1)$ for each strand, so that the set of endpoints of strands is some permutation of the set of origins. Two braids are equivalent if there is an orientation-preserving homeomorphism from $D^{2}\times I$ to $D^{2}\times I$ such that one braid is the image of the other and at each time during the homeomorphism the image is also a braid. For each $u\in\mathbb{S}^{1}$, the projection of a braid in the direction $u$ provides a braid diagram of $B$ with crossings. In this paper we will assume braid diagrams with generic crossings, and we further use that each braid can be represented in a piecewise vertical diagram, that is the strands in the diagram are vertical expect in a (small) neighbourhood of a crossing. This result seems to be folklore; see for example “Reidemeister’s Theorem" and the discussion in [4].
@@
A closed braid or braided link is the image of a braid under the map from the solid cylinder to the solid torus, $D^{2}\times I\to D^{2}\times\mathbb{S}^{1}{\subset\mathbb{R}^{2}\times\mathbb{C}\simeq\mathbb{R}^{4}}$, which sends $(x,y)\mapsto(x,e^{iy})$, where $e^{iy}$ gives the standard embedding from $\mathbb{R}/2\pi\mathbb{Z}$ into $\mathbb{C}\simeq\mathbb{R}^{2}$. Under the standard embedding $\mathcal{T}$ of the solid torus in $3$-dimensional Euclidean space, that is rotationally symmetric around the $z$-axis and orienting each strand positively the closed braid can be considered a link with $n\leq m$ components. Importantly, the orientation of each resulting component is aligned to a positive orientation on the core circle of the solid torus at all points. The braid index is the minimum number of strands required to form a closed braid equivalent to a given link.
@@
Of course this correspondence is not bijective as each link may be equivalent to many closed braids. An algorithmic alternative proof of this result was later given in [29]. They also showed that the number of elementary operations to obtain a braid diagram from a given link diagram with $n$ crossings and $p$ Seifert circles is at most $(p-1)(p-2)/2$ and the number of crossings in the resulting braid is at most $n+(p-1)(p-2)$. The braid index of a link is the smallest number of strands needed for a closed braid representation of the link. The braid index is equal to the minimal number of Seifert circles in any diagram of the braid [30]. We’ll now define the Seifert circles.
@@
Given an oriented link diagram, by eliminating each crossing and connecting incoming strands with adjacent outgoing strands we obtain a diagram of oriented circles known as Seifert circles.
2.3 Vineyards
We make the following blanket assumption: We assume that all our persistence diagrams only contain a finite number of points. We note that this is always true for $C^{2}$ manifolds with a tame Morse functions, which is the setting we require. [^1] Note that we are not referred to tame knots but rather tame Morse functions in this work, as an embedded tame knot may have an infinite number of critical points in its distance or height filtration.

### v2

@@
Braiding Vineyards
Erin W. Chambers Department of Computer Science and Engineering, University of Notre Dame, Notre Dame, IN (echambe2@nd.edu, https://wolfchambers.github.io/).    Christopher Fillmore Institute of Science and Technology Austria, Klosterneuburg, Austria (cdfillmore@gmail.com, https://orcid.org/0000-0001-7631-2885).    Elizabeth Stephenson Orteliu, Oslo, Norway (elizasteprene@gmail.com, https://orcid.org/0000-0002-6862-208X).    Mathijs Wintraecken Inria Centre Université Côte d’Azur, Sophia-Antipolis, France (mathijs.wintraecken@inria.fr, https://orcid.org/0000-0002-7472-2220).
In this work, we introduce and study what we believe is an intriguing and, to the best of our knowledge, previously unknown connection between two fundamental areas in computational topology, namely topological data analysis (TDA) and knot theory. Given a function from a topological space to $\mathbb{R}$, TDA provides tools to simplify and study the importance of topological features: in particular, the $l^{th}$-dimensional persistence diagram encodes the topological changes (or $l$-homology) in the sublevel set as the function value increases into a set of points in the plane. Given a continuous one-parameter family of such functions, we can combine the persistence diagrams into an object known as a vineyard, which track the evolution of points in the persistence diagram as the function changes. If we further restrict that family of functions to be periodic, we identify the two ends of the vineyard, yielding a closed vineyard. This allows the study of monodromy, which in this context means that following the family of functions for a period permutes the set of points in a non-trivial way. Recent work has studied monodromy in the directional persistent homology transform, demonstrating some interesting connections between an input shape and monodromy in the persistent homology transform for 0-dimensional homology embedded in $\mathbb{R}^{2}$.
In this work, given a link and a value $l$, we construct a topological space (based on the given link) and periodic family of functions on this space (based on the Euclidean distance function), such that the closed $l$-vineyard contains this link. This shows that vineyards are topologically as rich as one could possibly hope, suggesting many future directions of work. Importantly, it has at least two immediate consequences we explicitly point out:
  • Monodromy of any periodicity can occur in a $l$-vineyard for any $l$. This answers a variant of a question by Arya and collaborators [4]. To exhibit this as a consequence of our first main result we also reformulate monodromy in a more geometric way, which may be of interest in itself.
  • Topologically distinguishing closed vineyards is likely to be difficult (from a complexity theory as well as a practical perspective) because of the difficulty of knot and link recognition, which have strong connections to many NP-hard problems.
@@ 1 Introduction
Computational topology has a number of different branches which, despite some overlap, have remained fairly distinct in nature. On the one hand there is persistent homology [24, 41, 22], a more recent and active area which has its roots in algebraic topology; see [39, 20] for recent survey books on this active topic. On the other hand, there are the computational aspects of knot theory and the study and characterization of low dimensional manifold topology, which has a long history of algorithmic development, perhaps dating back to Dehn’s algorithm [18, 19] and the many following algorithmic results in more recent decades on shape and knot recognition in low dimensions [42, 46, 25, 6, 7].
@@
The first goal of this paper is to study an interesting new link between these two important branches of computational topology. Thanks to the work of Alexander [2] we know that every knot or link can be represented as a braid, that is, for every link there is a braid such that if we glue the ends of the braid, we recover the link; see Section 2.2 for more formal definitions. Recall also that for a continuous one parameter family of filtrations, we can “stack” the persistence diagrams of these filtrations; we call the resulting object a vineyard [15, 47, 27]. Thanks to the stability of persistence diagrams, the points in the persistence diagram move continuously (even Lipschitz continuously) with the parameter. This means that we can follow a point in (the stack of) the persistence diagrams; the resulting curve is called a vine. We will prove that for every link there exists an embedded manifold and a family of functions on $\mathcal{M}$ (where each function is induced by the distance to a point in the ambient space, and where in turn each point comes from a curve $\gamma$) such that the vineyard of the family of functions yields the braid representing the link in the sense of Alexander.
@@
The second goal of this paper is to show that any type of monodromy can occur in vineyards. This is part of a new research direction in computational topology. In [4], the occurrence of monodromy in the context of the directional persistence transform is studied, more precisely for 0-dimensional persistence modules of objects embedded in $\mathbb{\mathbb{R}}^{2}$. Roughly speaking, the directional persistence transform considers the persistence diagram of the height function on a shape for any direction, which in two dimensions gives what we call a closed vineyard (whose precise definition will be given below). It should be noted that monodromy had already been identified much earlier in the context of multidimensional persistence in [9], see also [37]. The authors of [4] conclude with an open question about demonstrating monodromy in higher dimensions, as well as several interesting and more open ended questions related to better understanding what monodromy is capturing about the input shape.
The setting of [4] is the directional persistence transform, however this has been placed in a larger context [38], where one studies the distance to flats of any dimension, including points. In this paper, we focus on that extreme case, namely points. The resulting radial transform allows for a more geometric understanding of monodromy, and is critical in our construction.
@@
In this paper, we exhibit that any type of monodromy and the braid associated to any link can occur in a vineyard. To make our statement more precise, however, we need to introduce some nomenclature, although full definitions will be deferred until Section 2. Intuitively, monodromy is the effect where if one makes a loop in a base space of a covering or fibre bundle, the lifted curve may not end up at the same point as you started out with. We say that the monodromy is of period $2\pi k$ (with $k>0$) if the lifted curve returns to the starting point after $k$ revolutions in the base space. In our context the base space is a closed curve or loop $\gamma:[0,2\pi]\to\mathbb{R}^{d}$, it is into this image $\mathbb{R}^{d}$ that we have embedded a manifold $\mathcal{M}$ (link or some offset of link, which is a modification of the input link). The fibres are the persistence diagrams of the Euclidean distance function restricted to the manifold $\mathcal{M}$, that is $d_{\mathbb{E}}(x,\gamma(t))|_{\mathcal{M}}$. The bundle therefore is the vineyard. The lifted curve is a vine $\tilde{\gamma}_{(b_{0},d_{0})}(t)$ in the vineyard starting at $(b_{0},d_{0})$ in the persistence diagram of $d(x,\gamma(0))_{\mathcal{M}}$ and the periodicity is the smallest $k>0$, such that for all $i$, $\tilde{\gamma}_{(b_{0},d_{0})}(0)=\tilde{\gamma}_{(b_{i},d_{i})}(2\pi k)$, where we assume that the vine is non-degenerate in the sense that it stays away from the diagonal.
@@
Theorem 1.1 (Any link occurs in some closed vineyard).
Given a link $\mathcal{X}$, $d,l\in\mathbb{Z}_{\geq 0}$, with $d\geq 3$ and l<d2l<d-2, then there exists a submanifold $\mathcal{M}\subset\mathbb{R}^{d}$ and a closed curve $\gamma\subset\mathbb{R}^{d}$ such that identifying the ends of the $l$-vineyard of $d(x,\gamma(t))|_{\mathcal{M}}$ yields a link which contains the given link as a subset. That is, the vineyard is topologically equivalent to the link we were given after removing some spurious connected components, i.e. the output is a link $\mathcal{Z}=\tilde{\mathcal{X}}\sqcup\mathcal{Y}$, where $\sqcup$ denotes the disjoint union of different connected components and $\tilde{\mathcal{X}}$ is ambiently isotopic to $\mathcal{X}$.
Theorem 1.2 (Any monodromy occurs in some closed vineyard).
The radial transform in $\mathbb{R}^{d}$ can exhibit monodromy for persistence up to the (d2d-2)th homology and for extended persistence up to the (d1d-1)th homology. Moreover the periodicity of the monodromy can be $2k\pi$ for any $k\in\mathbb{Z}_{\geq 2}$.
Figure 1.1: For a given a link (top left), we construct a closed braid (that is, a braid with ends identified) embedded near an annulus (bottom left). Note that we add an additional trivial loop in each connected braid component traveling around the exterior of the annulus. We then partially twist the annulus (top right) such that the cylindrical vertical portion in gray and the horizontal portion in yellow are perpendicular to each other. In the top right figure, the top edge is shown in red and the bottom edge in dark blue to clarify this twisting. It looks pinched due to the perspective, but is in fact only twisted. The bottom right figure shows the result of twisting the braid together with the annulus, with some massaging to keep the crossings in the horizontal part. Due to the addition of the exterior trivial loops from the bottom left figure, the vineyard consisting of the persistence diagrams of the distance function to a point following the twisted annulus (bottom right) on the ‘outside’ contains the input braid with some surgery. See Figure 1.2 for several 3D views of the embedded knot, as well as Figure A.2 and the corresponding detailed example and discussion in Appendix A.
Figure 1.2: Top: a side view of a 3-dimensional embedding of the link depicted in Figure 1.1, where the crossings and strands are perturbed in a particular way to get the desired monodromy (see the description in Figure 1.1). The red and gold portions are the two sections of the link, and the blue curve is an observation loop for the radial transform, where the dark blue portion corresponds to the section of vineyard visualized in the bottom. See also the proof of Theorem 1.2 and the corresponding Example in Appendix A, for full details on the perturbation. Middle: Front-angled view of the embedded link. Bottom: A sideways view of our computed vineyard, showing the diagrams computed from the radial transform of observation points taken from the fraction of the full period (from 0 to $2\pi$) depicted in dark blue segment of the curve above. This segment of the vineyard captures all crossings and exhibits monodromy of period $2\pi\cdot 3$. To transform this braided vineyard into a closed braid, we would identify the “sides”, or slices, at 0 and $2\pi$.
We remark that our construction hinges not only on $\mathcal{M}$ but also on a careful choice of $\gamma$. Indeed, fixing $\mathcal{M}$, there can be closed curves on $\mathcal{M}$ which yield no links in the vineyard.
@@ 2.1 Monodromy
Monodromy is an important concept in mathematics that appears in various guises. We refer to the review [21] (the first part of which is almost a review of reviews) and the other reviews mentioned in that paper for an overview of the various aspects of the theory. In this paper we will only consider the simplest incarnation, and only in the setting of topological data analysis.
Let $\tilde{X}$ be a covering space of $X$ with covering map $C:\tilde{X}\to X$, that is for every $x\in X$ there exists an open neighbourhood $x\in U$ and a discrete set $J$, such that $C^{-1}(U)=\sqcup_{i\in J}V_{i}$ and $C|_{V_{j}}:V_{j}\to U$ is a homeomorphism for all $j\in J$. See Figure 2.1 for an illustration. We call the inverse images of points $x\in X$ of the map $C$ the fibres. For a curve $\gamma:[0,2\pi]\to X$ we write $\tilde{\gamma}$ for (one of) its lift(s), that is a continuous map $\tilde{\gamma}:[0,2\pi]\to\tilde{X}$ such that $C\cdot\tilde{\gamma}=\gamma$. If $\gamma$ is a loop, that is $\gamma(0)=\gamma(2\pi)$, then we say that $\gamma$ exhibits monodromy (at the starting point $\tilde{\gamma}(0)\in C^{-1}(\gamma(0))$) if we have that the start and end points of its lift $\tilde{\gamma}$ are different, i.e. $\tilde{\gamma}(0)\neq\tilde{\gamma}(2\pi)$. The difference between $\tilde{\gamma}(0)$ and $\tilde{\gamma}(2\pi)$ is also referred to as monodromy (this difference can in certain cases be best represented by a group, see [21], although we will not need this in our discussion).
Figure 2.1: Here we see a cover $\tilde{X}$ (in this case a double cover) of the base space $X$, in this case a circle as well as the curve $\gamma$ and its lift $\tilde{\gamma}$.
If $\gamma$ is a loop we can extend it, formally speaking by concatenating with itself. Here we adopt the convention that if $\gamma$ and $\tilde{\gamma}$ are two curves parametrized by $0,2\pi]$, then the concatenation $\gamma\circ\tilde{\gamma}$ is parametrized by $0,4\pi]$, that is, we do not rescale the parametrization interval. We write
@@
where $\tilde{\gamma}$ is the lifting of $\gamma$. In fact the right hand side does not even have to be a continuous curve. We say that a loop $\gamma$ (parametrized by $0,2\pi]$) in the base space $X$ exhibits monodromy of order $k$ if $k$ is the smallest positive integer such that the lifted curve $\widetilde{\gamma^{k}}$ satisfies
@@ 2.2 Knots, links, and braids
In this subsection we briefly recall the formal definitions of knots, links and braids. An (oriented) knot is the the equivalence class of oriented closed curves embedded in $3$-dimensional Euclidean space, $\gamma\colon\mathbb{S}^{1}\to\mathbb{R}^{3}$, under ambient isotopy. To simplify the discussion, we will abuse notation and write $\gamma$ for the map as well as its image in $\mathbb{R}^{3}$. A link with $n$ components is a disjoint union of $n$ knots, $L=\gamma_{1},\cup\ldots\cup\gamma_{n}\subset\mathbb{R}^{3}$. We say two knots (or links) are equivalent if there is an orientation-preserving homeomorphism from $\mathbb{R}^{3}$ to $\mathbb{R}^{3}$ such that one knot (or link) is the image of the other. Generally, this equivalence is formalized as a series of Reidemeister moves (see Figure 2.2), as any isotopy between knots can be related by a sequence of these three local moves [3, 40].
Figure 2.2: The Reidemeister moves. Left to right: Type I, Type II, Type III.
For each $u\in\mathbb{S}^{2}$, the projection of a knot (or link), $\gamma$, in the direction $u$ provides a knot diagram (or link diagram) of $\gamma$ with crossings. In this paper we will assume diagrams with generic crossings, so that our diagram has a finite number of crossings and no self-tangencies.
A braid on $m$ strands is the disjoint union of $m$ intervals embedded in a solid [^1] That is the product of the disk $D^{2}$ and the interval $I$. cylinder, that is the braid $B$ is the union of strands $B_{i}$, where $B_{i}\colon I\to D^{2}\times I$, monotonically increasing with respect to $I$, such that $B_{i}(0)=(d_{i},0)$ and $B_{i}(1)=(d_{j},1)$ for each $i$, where $i$ indexes the strand, so that the set of endpoints of strands is some permutation of the set of origins. Two braids are equivalent if there is an orientation-preserving homeomorphism from $D^{2}\times I$ to $D^{2}\times I$ such that one braid is the image of the other and at each time during the homeomorphism the image is also a braid. For each $u\in\mathbb{S}^{1}$, the projection of a braid in the direction $u$ provides a braid diagram of $B$ with crossings. In this paper we will assume braid diagrams with generic crossings, and we further use that each braid can be represented in a piecewise vertical diagram, that is the strands in the diagram are vertical except in a (small) neighbourhood of a crossing. The fact that braids can always be represented in such a way seems to be folklore; see for example “Reidemeister’s Theorem" and the discussion in [5].
@@
A closed braid or braided link is the image of a braid under the map from the solid cylinder to the solid torus, $D^{2}\times I\to D^{2}\times\mathbb{S}^{1}{\subset\mathbb{R}^{2}\times\mathbb{C}\simeq\mathbb{R}^{4}}$, which sends $(x,y)\mapsto(x,e^{iy})$, where $e^{iy}$ gives the standard embedding from $\mathbb{R}/2\pi\mathbb{Z}$ into $\mathbb{C}\simeq\mathbb{R}^{2}$. The braid index is the minimum number of strands required to form a closed braid equivalent to a given link.
Under the standard embedding $\mathcal{T}$ of the solid torus in $3$-dimensional Euclidean space, where the embedding is rotationally symmetric around the $z$-axis and where each strand is oriented positively, the closed braid can be considered a link with $n\leq m$ components. Importantly, the orientation of each resulting component is aligned to a positive orientation on the core circle of the solid torus at all points.
@@
Of course this correspondence is not bijective as each link may be equivalent to many closed braids. An algorithmic alternative proof of this result was later given in [48]. The complexity of the algorithm depends on the number of Seifert circles, which are defined as:
@@
Given an oriented link diagram, by eliminating each crossing and connecting each incoming strand with its adjacent outgoing strand we obtain a diagram of oriented circles known as Seifert circles.
The number of elementary operations of the algorithm in [48] to obtain a braid diagram from a given link diagram with $n$ crossings and $p$ Seifert circles is at most (p1)(p2)/2(p-1)(p-2)/2 and the number of crossings in the resulting braid is at most n+(p1)(p2)n+(p-1)(p-2). The braid index of a link is the smallest number of strands needed for a closed braid representation of the link. The braid index is equal to the minimal number of Seifert circles in any diagram of the braid [51].
2.3 Persistence and Vineyards
We assume that the reader is familiar with the basic topics in algebraic topology [35, 26], Morse theory and handle decompositions [32], and the theory of persistent homology [20, 39], given the tools needed in our construction. We nevertheless give a very brief schematic overview of persistence because we need to contrast this with extended persistence, which is in the opinion of the authors not as utilized as it should be, and therefore may deserve to be recalled. However, we do not aim to be complete in our schematic overview, and refer the curious to [20, 39] for further details.
We fix the following blanket assumption throughout the entirety of the paper: We assume that all our persistence diagrams only contain a finite number of points (counted with multiplicity). We note that this is always true for $C^{2}$ manifolds with a tame Morse function, which is the setting we require. [^2] Note that we are not referring to tame knots but rather tame Morse functions in this work, as an embedded tame knot may have an infinite number of critical points in its distance or height filtration.